# ShareLoRA: Parameter Efficient and Robust Large Language Model Fine-tuning via Shared Low-Rank Adaptation

### v1

@@ Abstract
This study introduces an approach to optimize Parameter Efficient Fine Tuning (PEFT) for Pretrained Language Models (PLMs) by implementing a Shared Low Rank Adaptation (ShareLoRA). By strategically deploying ShareLoRA across different layers and adapting it for the Query, Key, and Value components of self-attention layers, we achieve a substantial reduction in the number of training parameters and memory usage. Importantly, ShareLoRA not only maintains model performance but also exhibits robustness in both classification and generation tasks across a variety of models, including RoBERTa, GPT-2, LLaMA and LLaMA2. It demonstrates superior transfer learning capabilities compared to standard LoRA applications and mitigates overfitting by sharing weights across layers. Our findings affirm that ShareLoRA effectively boosts parameter efficiency while ensuring scalable and high-quality performance across different language model architectures.
@@ 1 Introduction
* Equally contributed.
As Pretrained Language Models (PLMs) have gained prominence Devlin et al. (2019); Liu et al. (2019); Radford et al. (2019); Raffel et al. (2020), researchers are increasingly focused on optimizing the utilization of these models’ pre-trained weights. Traditional fine-tuning, which involves adjusting all parameters of a PLM for a specific dataset or task, is often resource-intensive and time-consuming, especially given the massive scale of large language models (LLMs) Brown et al. (2020); Kaplan et al. (2020); Hoffmann et al. (2022); Chowdhery et al. (2022); Zhang et al. (2022); Touvron et al. (2023a).
@@
Given the impressive performance of LoRA, numerous subsequent studies have aimed to enhance its efficiency, mainly by reducing the number of trainable parameters to minimize the memory footprint during the fine-tuning process. However, significantly lowering the trainable parameters can lead to slow convergence, while insufficient reductions may encourage the model to easily overfit. Therefore, we pose the question: Is there a PEFT approach that effectively balances trainable parameter selection, minimizes the memory footprint required for model parameters, and maintains the model’s adaptability?
To address this issue, we introduce ShareLoRA, an efficient and straightforward PEFT method that effectively balances trainable parameter selection while optimizing the model’s adaptability and minimizing memory requirements. Our approach leverages the observation that low-rank weight matrices A and B do not need to be uniquely configured across layers to achieve optimal PEFT performance in PLMs. Instead, we propose sharing either matrix A or B across all layers while maintaining its counterpart as distinct in each layer. This strategy meets several key objectives: 1) Sharing a low-rank matrix across layers significantly reduces the number of trainable parameters and cuts down on the memory footprint needed for model finetuning; 2) Keeping the shared matrix trainable preserves the model’s adaptability; 3) The updated weights for each component that LoRA applies remain unique yet share a common base.
In our experiments, we demonstrate the benefits of ShareLoRA under three configurations: 1) sharing across all layers, and 2) sharing the Query, Key, and Value components of the self-attention layers in PLMs. 3) sharing the down-projection, up-projection, or both in LoRA. The results show that ShareLoRA not only preserves model performance but also shows robustness in a variety of tasks, both in classification and generation, across multiple models including RoBERTa, GPT-2, and LLaMA. This method exhibits enhanced transfer learning capabilities compared to traditional LoRA applications and effectively prevents overfitting by sharing weights across layers. Our findings prove that ShareLoRA significantly improves parameter efficiency while maintaining scalable and high-quality performance across diverse language model architectures.
@@ 2 Related Work
Parameter Efficient Fine-tuning. PLMs are trained on large datasets to develop broad linguistic representations Devlin et al. (2019); Liu et al. (2019); Raffel et al. (2020), but often fall short in specialized tasks due to a lack of domain knowledge. Traditional approaches involve fully fine-tuning PLMs to enhance domain-specific performance Xu and Wang (2023); Xie et al. (2020); Dabre et al. (2019). However, with the increasing size of PLMs Workshop et al. (2023); Touvron et al. (2023a, b); Zhang et al. (2022), this method becomes too resource-heavy. As an alternative, Parameter Efficient Fine-tuning (PEFT) provides an efficient way to maintain performance with less computational expense.
@@
PEFT methods have become crucial for adapting large-scale pre-trained models to specific tasks without extensively overhauling their parameters. This approach conserves computational resources and boosts efficiency. For example, Prefix tuning Li and Liang (2021a) adds parameters to the hidden states across layers, subtly influencing the model’s behavior without changing its underlying architecture, Prompt tuning Lester et al. (2021) alters prompts and updates only the associated parameters, focusing on specific areas of model performance, and BitFit Zaken et al. (2022) updates only the biases within the model, resulting in minimal yet effective modifications.
@@
One notable PEFT technique is Low-Rank Adaptation (LoRA) Hu et al. (2021), which achieves efficient fine-tuning by incorporating a low-rank matrix adaptation mechanism alongside the existing weights of linear layers, thereby reducing memory overhead while preserving the effectiveness of the fine-tuning process. The modified output $Y$ is computed as follows:
$Y\leftarrow XW+\alpha XAB$ (1)
where $W$ represents the original pre-trained weights of dimensions $d_{\text{in}}\times d_{\text{out}}$, with $d_{\text{in}}$ being the dimension of the input to the layer, and $d_{\text{out}}$ being the dimension of the output. The input tensor $X$ has dimensions $b\times s\times d_{\text{in}}$ and the output tensor $Y$ has dimensions $b\times s\times d_{\text{out}}$, where $b$ and $s$ denote the batch size and sequence length, respectively.
The adaptation is facilitated by matrices $A$ and $B$, where $A\in\mathbb{R}^{d_{\text{in}}\times r}$ projects the input dimension down to a lower rank $r$, and $B\in\mathbb{R}^{r\times d_{\text{out}}}$ projects it back up, effectively creating a bottleneck that captures the most significant transformations. The hyperparameter $\alpha$, typically set inversely proportional to the rank $r$, scales the impact of this low-rank update on the output.
@@
Recent enhancements to LoRA have significantly broadened its capabilities. For instance, QLoRA Dettmers et al. (2023) optimizes LoRA for the fine-tuning of quantized models, thereby increasing efficiency. ReLoRA Lialin et al. (2023) incorporates a warm-up strategy during pre-training to boost adaptability. LoraHub Huang et al. (2024) streamlines the process by automating the creation of custom LoRA modules for specific tasks. Additionally, GLoRA Chavan et al. (2023) introduces a prompt module that fine-tunes weights and biases, enhancing performance across a variety of applications.
@@
Despite these advancements, LoRA still faces significant memory overhead due to the high activation memory usage in LoRA layers during the fine-tuning phase. To address this issue, LoRA-FA Zhang et al. (2023) strategically freezes the low-rank $A$ matrix and updates only the $B$ matrix. This approach significantly reduces the number of trainable parameters and activation memory, thus enhancing the efficiency of fine-tuning large language models without substantially impacting performance. However, LoRA-FA does not adequately decrease the total number of parameters that need to be stored, presenting a considerable challenge in contexts where computational resources and storage are constrained. Additionally, by freezing the $A$ matrix, LoRA-FA limits the model’s capacity to adapt and learn from new data during fine-tuning. This rigidity can hinder the model’s performance, particularly in complex or domain-specific tasks. Compared with LoRA-FA, our approach ShareLoRA offers a more dynamic and flexible strategy by allowing either matrix $A$ or $B$, or both, to be shared across different layers. This method not only preserves the model’s adaptability but also further reduces the memory requirements. We will show the details of it in the following paragraphs.
3 Approach
@@
In this section, we provide a detailed description of our proposed PEFT approach ShareLoRA, as illustrated in Figure1. ShareLoRA facilitates flexible configurations through two primary dimensions: 1) the choice of sharing between the matrices A, B, or both A and B (ShareA, ShareB, and ShareAB), and 2) the scope of sharing, which can be across different layers such as self-attention layers. This framework allows for a variety of combinations, enabling tailored adaptation of low-rank models to specific tasks. In the following paragraphs, we examine each configuration offered by ShareLoRA, exploring its benefits and implications in depth.
@@
ShareB Configuration
$\Delta W_{i}=\alpha A_{i}B=\alpha\sum_{k=1}^{r}A_{i,:,k}B_{k,:}$ (3)
ShareAB Configuration
$\Delta W=\alpha AB=\alpha\sum_{k=1}^{r}A_{:,k}B_{k,:}$ (4)
Sharing Across Self-Attention Layers
$Q_{i}$ $=X_{i}A_{Q}B_{Q_{i}}$ (5)
$K_{i}$ $=X_{i}A_{K}B_{K_{i}}$ (6)
$V_{i}$ $=X_{i}A_{V}B_{V_{i}}$ (7)
$\text{Attention}(Q_{i},K_{i},V_{i})$ $=\text{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{k}}}\right)V_{i},$ (8)
@@
4 Experiments
In our study, we conduct a comprehensive evaluation of the downstream performance of ShareLoRA across several series models, including RoBERTa Liu et al. (2019) and GPT-2 Radford et al. (2019). We benchmark these results against other established approaches such as LoRA Hu et al. (2021), LoRA-FA Zhang et al. (2023), on NLU and NLG tasks. Additionally, we extend the application of ShareLoRA to large-scale model in both LLaMA Touvron et al. (2023a) and LLaMA2 Touvron et al. (2023b) architectures, particularly in few-shot, zero-shot scenarios. Furthermore, our experiments cover a range of model sizes, from 7 billion to 13 billion parameters, and included both quantized and unquantized model variants. All tests were performed on the Nvidia A6000 and RTX 3090 GPUs.
4.1 Datasets
Additionally, we expand our experiments to few-shot and zero-shot tasks on larger models, demonstrating our approach’s adaptability. Following the configuration outlined in Dettmers et al. (2023), we employ Alpaca Taori et al. (2023) for LoRA and ShareLoRA, using the MMLU benchmark  Hendrycks et al. (2021) for evaluation. Some other benchmarks like ARC Chollet (2019), Hellaswrag Zellers et al. (2019) and GSM8K Cobbe et al. (2021) are used for comparison of model adaptability. All experimental setups are consistent with those described studies and demonstration of their repositories, based on the best of our knowledge.
4.2 Baselines
Full Fine-Tuning (FT) is a commonly used approach for model adaptation. It involves initializing the model with pre-trained weights and biases, and then gradient update all model’s parameters.
Bitfit is a technique studied by Zaken et al. (2022) for updating only a select small subset of biases parameters, to improve performance on new tasks while freezing all other pre-trained weights.
PreLayer Li and Liang (2021b) is a parameter-efficient technique for customizing large language models by learning specific activations after each Transformer layer for designated prefix tokens, while the main model parameters remain unchanged.
Adapter as introduced by Houlsby et al. (2019), involves inserting adapter layers between neural modules such as the self-attention and MLP modules, enhancing model flexibility without extensive modifications. AdapterL Lin et al. (2020) introduce adapters only after the MLP module followed by a LayerNorm, with AdapterD Rücklé et al. (2021) increases efficiency by omitting some adapter layers.
LoRA-FA Zhang et al. (2023) is a memory-efficient approach to fine-tuning large language models by reducing the activation memory required. This method innovatively freezes the projection-down weight of A and updates the up projection weight in each LoRA layer.
QLoRA Dettmers et al. (2023) utilizes a frozen, 4-bit quantized pretrained model and LoRA for efficient gradient propagation. The QLoRA approach features several innovations including 4-bit NormalFloat (NF4) data type and double quantization for memory reduction.
@@ 4.2 Baselines
GPT-2 M (ShareA) 0.20M $\mathbf{69.7_{\pm.4}}$ $8.75_{\pm.05}$ $\mathbf{46.60_{\pm.1}}$ $\mathbf{71.63_{\pm.1}}$ $\mathbf{2.51_{\pm.01}}$
@@
GPT-2 L (LoRA) 0.77M $69.8_{\pm.4}$ $8.80_{\pm.04}$ $\mathbf{46.69_{\pm.1}}$ $71.71_{\pm.3}$ $\mathbf{2.52_{\pm.01}}$
@@
5 Main Results
5.1 GLUE Benchmark
ShareA outperforms LoRA variants.
In Table1, we present the performance metrics for different versions of ShareLoRA—ShareA, ShareB, and ShareAB—alongside a baseline comparison with previously published work using RoBERTa-base and RoBERTa-large models.
For the RoBERTa-base model, ShareA demonstrates its strengths on datasets such as MRPC, CoLA, and RTE, where we notice performance improvements between 0.2% to 0.5%. This enhancement is noteworthy especially, under the same training specifications Hu et al. (2021), these datasets have reached full convergence and are prone to overfitting.
ShareA is adaptable and robust.
In tasks such as MRPC, RTE, and STS-B, both ShareLoRA and LoRA utilize the best MNLI checkpoint derived from multiple seeds and applies these checkpoints effectively on other tasks, demonstrating superior adaptability and performance enhancement compared to using LoRA alone once convergence is achieved. This adaptability highlights the potential of ShareLoRA in generalizing well across converged datasets.
ShareLoRA also has a marginal decline in performance as observed on the MNLI, QNLI, and QQP datasets compared to LoRA in Table1. Due to the large size of datasets, both LoRA and ShareLoRA are not fully converged under the configurations as described in Hu et al. (2021). However, it is crucial to highlight that even with the reduced performance on MNLI checkpoint, the adaptive tasks such as MRPC and RTE, still show better performance, underscoring the robustness of ShareLoRA, effectively preventing overfitting and optimizing performance outcomes.
ShareA outperforms ShareB.
Experiments conducted with the RoBERTa-large model on ShareA, ShareB, and ShareAB reveal that ShareA generally outperforms ShareB in various tasks and both ShareA and ShareB show superior results over ShareAB. Compared to LoRA, ShareA demonstrates increased stability with less fluctuation in the confidence intervals across the majority of tasks in Table1, emphasizing ShareLoRA’s advantage in providing consistent and reliable performance enhancements.
Parameter Efficiency of ShareLoRA
Additionally, our shared approach significantly reduces the number of trainable parameters compared to LoRA and other approaches. Employing a similar number of trainable parameters as LoRA-FA, but ShareLoRA achieves enhanced performance across all datasets.
Overall, the distinct advantages of ShareLoRA, particularly in terms of its efficiency, robustness, and adaptability to different NLU tasks leading to superior performance. ShareLoRA produces a compelling balance between performance and computational efficiency.
5.2 E2E Challenge
ShareA outperforms LoRA in NLG.
In Table2, we utilize the configuration previously outlined in Hu et al. (2021) with GPT-2 medium and large for E2E NLG tasks, showcasing the superiority of ShareLoRA in generative tasks. Our results indicate that ShareLoRA achieves a consistent performance improvement over LoRA across all evaluated metrics for the GPT-M model. When employing the GPT-large model, ShareLoRA demonstrates slightly better performance than LoRA, given that ShareLoRA utilizes only half the training parameters of LoRA, achieving a performance improvement of 0.1% to 0.2% over LoRA.
Up Projection Enhances Performance.
Furthermore, both LoRA and ShareA outperform ShareB in generative tasks across all metrics. Within the LoRA framework, the significance of the up-projection matrix B is evident as it crucially augments the dimensionality of the low-rank representation. The strategic choice to share component A rather than B in ShareLoRA proves advantageous, as it expansion the intermediate dimension is more important and difficult than squeezing the high dimension features in complex generation tasks.
@@ 5.2 E2E Challenge
5.3 LLaMA on MMLU
ShareA and ShareAqkv outperform LoRA.
In Table3, the scalability and efficacy of ShareA are assessed by examining its performance on larger models ranging from 7B to 13B parameters. Through fine-tuning on the Alpaca dataset and employing the 5-shot MMLU benchmark as specified by Dettmers et al. (2023), ShareA demonstrates notable enhancements in generative capabilities compared to GPT-2 and RoBERTa.Focusing exclusively on ShareA rather than ShareB, the results from different linear components indicate that LLaMA models, particularly the 13B and both the 7B and 13B versions of LLaMA2, outperform standard LoRA with improvements of approximately 1.1%, 0.7%, and 0.4%, respectively. Moreover, ShareAqkv further improves performance by 0.6% for the LLaMA 13B model over ShareA, while ShareA outperforms ShareAqkv by 0.3% for the LLaMA2 7B model. The closely matched performance between ShareAqkv and ShareA across other models suggests a high convergence and potential overfitting risks, as discussed in Appendix 5.3 and Figure4, with the LLaMA 7B model showing stable yet under-converged performance according to prior research Xu et al. (2023).
Memory Footprint Saving
In the context of smaller models like RoBERTa and GPT-2, ShareA yields minimal parameter savings, which is negligible given modern GPU capacities. However, with larger models like LLaMA, ShareA demonstrates more substantial reductions. Specifically, the LLaMA 7B and 13B models cut down approximately 60 million and 110 million trainable parameters, respectively, when compared to the LoRA architecture. This leads to substantial efficiency gains, reducing both computational footprint and disk storage needs. As depicted in Figure2 in the Appendix, ShareA achieves a memory footprint reduction of 1.8GB and approximately a 2% increase in training speed, while ShareAB can save around 4GB with similar training speeds. The confidence intervals in Table3 illustrate that ShareA not only improves performance but also increases robustness over standard LoRA, underscoring the practical advantages of ShareLoRA in LLMs.

### v2

@@ Abstract
In this paper, we introduce Shared Low Rank Adaptation (ShareLoRA), a Large Language Model (LLM) fine-tuning technique that balances parameter efficiency, adaptability, and robustness without compromising performance. By strategically sharing the low-rank weight matrices across different layers, ShareLoRA achieves 44% to 96% reduction in trainable parameters compared to standard LoRA, alongside a substantial decrease in memory overhead. This efficiency gain scales with model size, making ShareLoRA particularly advantageous for resource-constrained environments. Importantly, ShareLoRA not only maintains model performance but also exhibits robustness in both classification and generation tasks across diverse models, including RoBERTa, GPT-2, and LLaMA series (1, 2, and 3). It consistently outperforms LoRA in zero-shot, few-shot, and continual fine-tuning scenarios, achieving up to 1.2% average accuracy improvement, and enhanced generalization across domains. In continual learning settings, ShareLoRA achieves 1.2% higher accuracy on GSM8K, 0.6% on HumanEval, and 0.5% on both MMLU and MMLU-Pro. Our results demonstrate that ShareLoRA supports high-quality fine-tuning while offering strong generalization and continual adaptation across various model scales and diverse tasks. [^1] https://github.com/Rain9876/ShareLoRA
ShareLoRA: Parameter Efficient and Robust Large Language Model Fine-tuning via Shared Low-Rank Adaptation
Yurun Song UC Irvine yuruns@uci.edu                        Junchen Zhao UC Irvine junchez3@uci.edu
Ian G. Harris UC Irvine harris@ics.uci.edu                        Sangeetha Abdu Jyothi UC Irvine, VMware Research sangeetha.aj@uci.edu
@@ 1 Introduction
As Pretrained Language Models (PLMs) have gained prominence Devlin et al. (2019); Liu et al. (2019); Radford et al. (2019), researchers are increasingly focused on optimizing the utilization of these models’ pre-trained weights. Traditional fine-tuning, which involves adjusting all parameters of a PLM for a specific dataset or task, is often resource-intensive and time-consuming, especially given the massive scale of large language models (LLMs) Brown and et.al (2020); Kaplan et al. (2020); Hoffmann and et.al (2022); et.al (2022); Zhang et al. (2022); et.al (2023b).
@@
Given the impressive performance of LoRA, subsequent studies have aimed to enhance its efficiency, mainly by reducing the number of trainable parameters to minimize the memory footprint during the fine-tuning process. However, significantly lowering the trainable parameters can lead to slow convergence, while insufficient reductions may encourage the model to easily overfit. Moreover, existing PEFT methods often struggle to maintain robustness across different domains after fine-tuning.
To address these challenges, we introduce ShareLoRA, an efficient and straightforward PEFT method that effectively balances trainable parameter selection while optimizing the model’s adaptability, minimizing memory requirements, and ensuring robustness across domains. Our approach leverages the observation that low-rank weight matrices A and B do not need to be uniquely configured across layers to achieve optimal PEFT performance in PLMs. Instead, we propose sharing either matrix A or B across all layers while maintaining its counterpart as distinct in each layer. This strategy meets several key objectives:
  • Parameter Efficiency: Sharing a low-rank matrix across layers reduces trainable parameters by 44% to 96% compared to standard LoRA, for models such as LLaMA-7B. This memory reduction scales with model size which is critical for efficient fine-tuning LLMs on consumer GPUs and edge devices.
  • Model Adaptability: Keeping the shared matrix trainable preserves the model’s adaptability, allowing it to effectively learn and adapt to new tasks and domains. Also, the updated weights for each component that LoRA applies remain unique yet share a common base, promoting consistency across layers while allowing for task-specific adaptations.
  • Continual Adaption: ShareLoRA exhibits robustness when continual fin-tuning to domains different from the one it was fine-tuned on. This generalization capability sets it apart from traditional LoRA and other PEFT methods, which often struggle to maintain performance when faced with out-of-domain tasks.
Our extensive experiments across multiple models, including RoBERTa, GPT-2, and LLaMA series, demonstrate that ShareLoRA not only preserves model performance but also shows remarkable robustness across a variety of tasks in both classification and generation.
@@ 2 Related Works
PLMs are trained on large datasets to develop broad linguistic representations Devlin et al. (2019); Liu et al. (2019); Raffel et al. (2020), but often fall short in specialized tasks due to a lack of domain knowledge. Traditional approaches involve fully fine-tuning PLMs to enhance domain-specific performance Xu and Wang (2023); Xie et al. (2020); Dabre et al. (2019). However, with the increasing size of PLMs Workshop et al. (2023); et.al (2023b, a); Zhang et al. (2022), this method becomes too resource-heavy. As an alternative, Parameter Efficient Fine-tuning (PEFT) provides an efficient way to maintain performance with less computational expense.
@@
PEFT methods have become crucial for adapting large-scale pre-trained models to specific tasks without extensively overhauling their parameters. This approach conserves computational resources and boosts efficiency. For example, Prefix tuning Li and Liang (2021a) adds parameters to the hidden states across layers, subtly influencing the model’s behavior without changing its underlying architecture. Prompt tuning Lester et al. (2021) alters prompts and updates only the associated parameters, focusing on specific areas of model performance. BitFit Zaken et al. (2022) updates only the biases within the model, resulting in minimal yet effective modifications.
@@
One notable PEFT technique is Low-Rank Adaptation (LoRA) Hu et al. (2021), which achieves efficient fine-tuning by incorporating a low-rank matrix adaptation mechanism alongside the existing weights of linear layers. This approach reduces memory overhead while preserving the effectiveness of the fine-tuning process.
@@
Recent enhancements to LoRA have significantly broadened its capabilities. QLoRA Dettmers et al. (2023) optimizes LoRA for the fine-tuning of quantized models, thereby increasing efficiency. ReLoRA Lialin et al. (2023) incorporates a warm-up strategy during pre-training to boost adaptability. LoraHub Huang et al. (2024) streamlines the process by automating the creation of custom LoRA modules for specific tasks. Additionally, GLoRA Chavan et al. (2023) introduces a prompt module that fine-tunes weights and biases, enhancing performance across a variety of applications.
@@
Despite these advancements, LoRA still faces significant memory overhead due to high activation memory usage in LoRA layers during the fine-tuning phase. To address this issue, LoRA-FA Zhang et al. (2023) strategically freezes the low-rank $A$ matrix and updates only the $B$ matrix. This approach significantly reduces the number of trainable parameters and activation memory, thus enhancing the efficiency of fine-tuning large language models without substantially impacting performance.
However, LoRA-FA does not adequately decrease the total number of parameters that need to be stored, presenting a considerable challenge in contexts where computational resources and storage are constrained. Additionally, by freezing the $A$ matrix, LoRA-FA limits the model’s capacity to adapt and learn from new data during fine-tuning. This rigidity can hinder the model’s performance, particularly in complex or domain-specific tasks.
In contrast, our proposed approach ShareLoRA offers a more dynamic and flexible strategy by allowing either matrix $A$ or $B$, or both, to be shared across different layers. This method not only preserves the model’s adaptability but also further reduces the memory requirements.
3 Method
@@
In this section, we provide a detailed description of our proposed PEFT approach ShareLoRA, as illustrated in Figure 1. ShareLoRA facilitates flexible configurations through two primary dimensions: 1. the choice of sharing between the matrices A, B, or both A and B (ShareA, ShareB, and ShareAB), and 2. the scope of sharing, which can be across different layers such as self-attention layers. This framework allows for a variety of combinations, enabling tailored adaptation of low-rank models to specific tasks.
@@
[…]
ShareB Configuration
$\Delta W_{i}=\alpha A_{i}B=\alpha\sum_{k=1}^{r}A_{i,:,k}B_{k,:}$ (2)
ShareAB Configuration
$\Delta W=\alpha AB=\alpha\sum_{k=1}^{r}A_{:,k}B_{k,:}$ (3)
Sharing Across Self-Attention Layers
$Q_{i}$ $=X_{i}A_{Q}B_{Q_{i}}$ (4)
$K_{i}$ $=X_{i}A_{K}B_{K_{i}}$ (5)
$V_{i}$ $=X_{i}A_{V}B_{V_{i}}$ (6)
$\text{Attention}(Q_{i},K_{i},V_{i})$ $=\text{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{K_{i}}}}\right)V_{i},$ (7)
@@
GPT-2 M (ShareA) 0.20M $69.7_{\pm.4}$ $8.75_{\pm.05}$ $\mathbf{46.60_{\pm.1}}$ $\mathbf{71.63_{\pm.1}}$ $\mathbf{2.51_{\pm.01}}$
@@
GPT-2 L (VERA)* 0.17M 70.3 8.85 46.9 $71.6$ $2.52$
GPT-2 L (LoRA) 0.77M $69.8_{\pm.4}$ $8.80_{\pm.04}$ $46.69_{\pm.1}$ $71.71_{\pm.3}$ $\mathbf{2.52_{\pm.01}}$
@@
4 Experiments
In our study, we conduct a comprehensive evaluation of the downstream performance of ShareLoRA across several series models, including RoBERTa Liu et al. (2019) and GPT-2 Radford et al. (2019). We benchmark these results against other established approaches such as LoRA Hu et al. (2021), LoRA-FA Zhang et al. (2023). Additionally, we extend the application of ShareLoRA to large-scale model in LLaMA series (et.al, 2023b, et.al, 2023a, Dubey et al., 2024) architectures, particularly in few-shot, zero-shot scenarios. Furthermore, our experiments cover a range of model sizes, from 7 billion to 13 billion parameters, and included both quantized and unquantized model variants. All tests were performed on the Nvidia A6000 and RTX 3090 GPUs. For experiment hyper-parameter settings, see Appendix Table 8-Table 11.
@@ 4 Experiments
4.1 Datasets
Additionally, we expand our experiments to few-shot and zero-shot tasks on larger models, demonstrating our approach’s adaptability. Following the configuration outlined in Dettmers et al. (2023), we employ Alpaca Taori et al. (2023), CodeAlpacaChaudhary (2023) and MATH Hendrycks et al. (2021b) for LoRA and ShareLoRA, using the MMLU benchmark  Hendrycks et al. (2021a) for evaluation. Some other benchmarks like ARC Chollet (2019), Hellaswrag Zellers et al. (2019), MMLU-Pro Wang et al. (2024), HumanEvalChen et al. (2021) and GSM8K Cobbe et al. (2021) are used for comparison of model adaptability. All experimental setups are consistent with those described studies and demonstration of their repositories, based on the best of our knowledge.
4.2 Baselines
Full Fine-Tuning (FT) is a commonly used approach for model adaptation involving with updating all model’s parameters.
Bitfit Zaken et al. (2022) is a technique for updating only a select small subset of biases parameters, to improve performance on new tasks while freezing all other pre-trained weights.
PreLayer/Prefix Li and Liang (2021b) is a parameter-efficient technique for customizing large language models by learning specific activations after each Transformer layer for designated prefix tokens, while the main model parameters remain unchanged.
Adapter Houlsby et al. (2019) involves inserting adapter layers between neural modules such as the self-attention and MLP modules, enhancing model flexibility without extensive modifications. AdapterL Lin et al. (2020) introduces adapters after the MLP module followed by a LayerNorm, while AdapterD Rücklé et al. (2021) increases efficiency by omitting some adapter layers.
IA3 Liu et al. (2022) is a PEFT approach that enhances model performance by scaling activations with learned vectors.
LoRA-FA Zhang et al. (2023) is a memory-efficient approach to fine-tuning large language models by reducing the activation memory required.
VERA Kopiczko et al. (2023) reduces trainable parameters by using frozen random matrices and learned scaling vectors, matching LoRA’s performance more efficiently.
Tied-LoRA Renduchintala et al. (2023) improves parameter efficiency by tying weights and training fewer low-rank matrices, matching LoRA performance with significantly fewer parameters.
VB-LoRA Li et al. (2024) achieves extreme parameter efficiency by generating low-rank adaptation weights from a shared vector bank using a differentiable top-k selection.